# Higher dimensional worm domains

### v1

@@ 1. Introduction
    the restriction of its D’Angelo class to $Y$ is a prescribed element of $H^{1}_{\mathrm{dR}}(Y,\mathbb{R})$,
@@
    every boundary point not lying on the closure of $Y$ is strongly pseudoconvex.
@@
Let $X$ be a Stein manifold. Let $Y\subset X$ be a precompact domain with strongly pseudoconvex boundary of class $C^{k}$ ($k\geq 2$, possibly $\infty$). Let $\gamma\in H^{1}_{\mathrm{dR}}(Y,\mathbb{R})$. Let $d\geq 1$ be an integer. Then there exists a domain $W$ in $X\times\mathbb{C}^{d}$ having the following properties:
@@
A three-dimensional variant of the Diederich–Fornaess worm domain is the object of the recent preprint [21]. The boundary of the domain considered by these authors contains the product of two annuli and is weakly pseudoconvex of Levi rank one on an open subset (see the last paragraph of Section 2 for a comment related to this point). Notice that the methods discussed below cannot be directly applied to construct a domain which is strongly pseudoconvex outside a bi-annulus, as this is not strongly pseudoconvex (and only Lipschitz regular). It would certainly be of interest to generalize Theorem 1 by relaxing the strong pseudoconvexity (and the $C^{2}$ regularity) condition. Notice that the hypothesis that $Y$ be weakly pseudoconvex cannot be dispensed with. See [6, Remark 2.2].
@@ 4. Proof of Theorem 1
It is also clear that 00 is a regular value of $R-\eta$ if and only if $K$ is a regular value of $e^{\frac{1}{d}}-\sigma$, viewed as a function on $X\setminus\overline{Y}$. By Sard Theorem, the set of regular values is dense, and the claim above follows. Thus, items (1) and (2) of Theorem 1 hold, for arbitrarily large choices of $K$. We now look at item (3).
@@ 5. Further properties and some open problems
Let $W$ be one of the domains constructed in Section 4. Denoting by $T^{1,0}bW$ the bundle of vectors of type $(1,0)$ tangent to the boundary, we may consider the Levi null distribution

### v2

@@ 1. Introduction
    every boundary point not lying on the closure of $Y$ is strongly pseudoconvex,
@@
    the restriction of its D’Angelo class to $Y$ is a prescribed element of $H^{1}_{\mathrm{dR}}(Y,\mathbb{R})$.
@@
Let $X$ be a Stein manifold. Let $Y\subset X$ be a precompact domain with strongly pseudoconvex boundary of class $C^{k}$, where $k\geq 2\dim_{\mathbb{C}}Y$ (possibly $\infty$). Let $\gamma\in H^{1}_{\mathrm{dR}}(Y,\mathbb{R})$. Let $d\geq 1$ be an integer. Then there exists a domain $W$ in $X\times\mathbb{C}^{d}$ having the following properties:
@@
A three-dimensional variant of the Diederich–Fornaess worm domain is the object of the recent preprint [21]. The boundary of the domain considered by these authors contains the product of two annuli and is weakly pseudoconvex of Levi rank one on an open subset (see the last paragraph of Section 2 for a comment related to this point). Notice that the methods discussed below cannot be directly applied to construct a domain which is strongly pseudoconvex outside a bi-annulus, as this is not strongly pseudoconvex (and only Lipschitz regular). It would certainly be of interest to generalize Theorem 1 by relaxing the strong pseudoconvexity (and the $C^{k}$ regularity) condition. Notice that the hypothesis that $Y$ be weakly pseudoconvex cannot be dispensed with. See [6, Remark 2.2].
@@ 4. Proof of Theorem 1
It is also clear that 00 is a regular value of $R-\eta$ if and only if $K$ is a regular value of $e^{\frac{1}{d}}-\sigma$, viewed as a $C^{k}$ function on $X\setminus\overline{Y}$. By Sard Theorem, which is applicable because $k$ is at least the real dimension of $X$, the set of regular values is dense, and the claim above follows. Thus, items (1) and (2) of Theorem 1 hold, for arbitrarily large choices of $K$. We now look at item (3).
@@ 5. Further properties and some open problems
Let $W$ be one of the domains constructed in Section 4. Assume that $W$ is $C^{\infty}$-smooth. Denoting by $T^{1,0}bW$ the bundle of vectors of type $(1,0)$ tangent to the boundary, we may consider the Levi null distribution